what is the solution to the system of equations of 2x-7y=17 and 3x+4y=11
Answer by mananth(16949)

2 x -7 y = 17 .............1
3 x + 4 y = 11 .............2
Eliminate y
multiply (1)by 4
Multiply (2) by 7
8 x -28 y = 68
21 x + 28 y = 77
Add the two equations
29 x = 145.00
/ 29
x = 5
plug value of x in (1)
2 x + -7 y = 17
10 + -7 y = 17
-7 y = 17 -10
-7 y = 7
y = -1
(5,-1)